New noise spectra of the green locking have been updated.
The plot below shows the in-loop noise spectra when the beat signal was fedback to ETMX.
The rms noise integrated from 0.1 Hz to 100 Hz went down to approximately 2 kHz.
The red curve was taken when the beat was controlled only by a combination of some poles sand zeros on the digital filter banks. The UGF was at 40Hz.
This curve is basically the same as that Koji took few weeks ago (see here). Apparently the rms was dominated by the peaks at 16 Hz and 3 Hz.
The blue curve was taken when the same filter plus two resonant gain filters (at 16.5 Hz and 3.15 Hz) were applied. The UGF was also at 40Hz.
Due to the resonant gain filter at 16.5 Hz, the phase margin became less, and it started oscillating at the UGF as shown in the plot.